# Matryoshka: Non-Exclusive Memory Tiering via Transactional Page Migration

### v1

@@ 2.3 Related Work
A long line of pioneering work has explored a wide range of of tiered storage/memory systems, built upon SSDs and HDDs [12, 15, 20, 54, 46, 38, 49, 30, 24], DRAM and disks [19, 23, 26, 43], HBM and DRAMs [42, 21, 45], NUMA memory [2, 3], PM and DRAM [37, 13, 40, 14], local and far memory [17, 25, 31, 32, 44], DRAM and CXL memory [35, 41, 34], and multiple tiers [48, 29, 51, 37, 33]. We focus on tiered memory systems consisting of DRAM and emerging CXL memory, though the proposed approaches in Matryoshka also apply to other tiered memory systems such as HBM/DRAM and DRAM/PM. We discuss the most related works as follows.
@@
Lightweight memory access tracking: To mitigate software overhead associated with memory access tracking, Hotbox [17] employs two separate scanners for fast and slow tiers to scan the slow tier at a fixed rate while the local tier at an adaptive rate, configurable based on the local memory pressure. Memtis [34] adjusts its PEBS-based sampling rate to ensure its overhead is under a threshold (e.g., < 3%). TMTS [22] also adopts a periodic scanning mechanism to detect frequency along with hardware sampling to more timely detect newly hot pages. While these approaches balance scanning/sampling overhead and tracking accuracy, an “always-on” profiling component does not seem practical, especially for high-pressure workloads. Instead, thermostat [14] samples a small fraction of pages, while DAMON [3] monitors memory access at a coarser-grained granularity (i.e., region). Although both can effectively reduce the scanning overhead, coarse granularity leads to lower accuracy regarding page access patterns. On the other hand, to reduce the overhead associated with frequent hint page faults like AutoNUMA [2], TPP [41] enables the page-fault based detection only for CXL memory (i.e., the slow tier) and tries to promote a page promptly (e.g., when it encounters a page fault); prompt page promotion avoids subsequent page faults on the same page. Inspired by existing lightweight tracking systems, such as Linux’s active and inactive lists and hint page faults, Matryoshka advances them by incorporating more recency information with no additional CPU overhead. Unlike hardware-assisted approaches [34, 22, 39], Matryoshka does not require any additional hardware support.
@@
Page migration optimizations: To hide reclamation overhead from applications, TPP [41] decouples page allocation and reclamation; however, page migration remains in the critical path, incurring significant slowdowns. Nimble [51] focuses on mitigating page migration overhead with new migration mechanisms, including transparent huge page migration and concurrent multi-page migration. Memtis [34] further moves page migration out of the critical path using a kernel thread to promote/demote pages in the background. TMTS [22] leverage a user/kernel collaborative approach to control page migration. In contrast, Matryoshka aims to achieve prompt, on-demand page migration while moving page migration off the critical path. It is orthogonal to and can benefit from existing page migration optimizations. The most related work is [18], which leverages hardware support to pin data in caches, enabling access to pages during migration. Again, Matryoshka does not need additional hardware support.
3 Matryoshka Design and Implementation
@@
Matryoshka is a new page management mechanism for tiered memory that features non-exclusive memory tiering and transactional page migration. The goal of Matryoshka design is to enable the processor to freely and directly access pages from both fast and slow memory tiers and move the cost of page migration off the critical path of users’ data access. Note that Matryoshka does not make page migration decisions and relies on memory tracing in the OS to determine hot and cold pages. Furthermore, Matryoshka does not impact the initial memory allocation in the OS and assumes a standard page placement policy. Pages are allocated from the fast tier whenever possible and are placed in the slower tier only when there is an insufficient number of free pages in the fast tier, and attempts to reclaim memory in the fast tier have failed. After the initial page placement, Matryoshka gradually migrates hot pages to the fast tier and old pages to the slow tier. Matryoshka seeks to address two key issues: 1) how to minimize the cost of page migration? 2) how to minimize the number of migrations?
@@
Overview. Inspired by multi-level cache management in modern processors, which do not employ a purely inclusive or exclusive caching policy between tiers [16] to facilitate the sharing of or avoid the eviction of certain cache lines, Matryoshka embraces a non-exclusive memory tiering policy to prevent memory thrashing when under memory pressure. Unlike the existing page management schemes that move pages between tiers and require that a page is only present in one tier, Matryoshka instead copies pages from the slow tier to the fast tier and keeps a shadow copy of the migrated pages at the slow tier. The non-exclusive tiering policy maintains shadow copies only for pages that have been promoted to the fast tier, thereby not an inclusive policy. The advantage of the non-exclusive policy is that the demotion of clean, cold pages can be simplified to remapping the page table entry (PTE) without the need to actually copy the cold page to the slower tier.
@@
The building block of Matryoshka is a new transactional page migration (TPM) mechanism to reduce the cost of page migrations. Unlike the existing unmap-copy-remap 3-step page migration, TPM opportunistically copies a page without unmapping it from the page table. During the page copy, the page is not locked and can be accessed by a user program. After the copy is completed, TPM checks if the page has been dirtied during the copy. If not, TPM locks the page and remaps it in the PTE to the faster tier. Otherwise, the migration is aborted and will be tried at a later time. TPM not only minimizes the duration during which a page is inaccessible but also makes page migration asynchronous, thereby removing it from the critical path of users’ data access.
@@
Without loss of generality, we describe Matryoshka design in the context of Linux. We start with transactional page migration and then delve into page shadowing – an essential mechanism that enables non-exclusive memory tiering.
@@ 3.1 Transactional Page Migration
The motivation to develop TPM is to make page migration entirely asynchronous and decoupled from users’ access to the page. As discussed in Section 2.2, the current page migration in Linux is synchronous and on the critical path of users’ data access. For example, the default tiered memory management in Linux, TPP, attempts to migrate a page from the slow tier whenever a user program accesses the page. Since the page is in “protected” mode, the access triggers a minor page fault, leading TPP to attempt the migration. The user program is blocked and makes no progress until the minor page fault is handled and the page is remapped to the fast tier, which can be a time-consuming process. Worse, if the migration fails, the OS remains in function migrate_pages and retries the aforementioned migration until it is successful or reaching a maximum of 10 attempts.
TPM decouples page migration from the critical path of user programs by making the migrating page accessible during the migration process. Therefore, users will access the migrating page from the slow tier before the migration is complete. While accessing a hot page from the slow tier may lead to sub-optimal memory performance, it avoids blocking user access due to the migration, thereby leading to superior user-perceived performance. Figure 3 shows the workflow of TPM. Before migration commences, TPM clears the protection bit of the page frame and adds the page to a migration pending queue. Since the page is no longer protected and it is not yet unmapped from the page table, following accesses to the page will not trigger additional page faults.
@@
Page migration is a complex procedure that involves memory tracing and updates to the page table for page remapping. The state-of-the-art page fault-based migration approaches, e.g., TPP in Linux [41], employs synchronous page migration, a mechanism in the Linux kernel for moving pages between NUMA nodes. In addition to the extended migration time affecting the critical path of user programs, this mechanism causes excessive page faults when integrated with the existing LRU-based memory tracing. TPP makes per-page migration decisions based on whether the page is on the active LRU list. Nevertheless, in Linux, memory tracing adds pages from the inactive to the active LRU list in batches of 15 requests  [^2] The 15 requests could be repeated requests for promoting the same page to the active LRU list, aiming to minimize the queue management overhead. Due to synchronous page migration, TPP may submit multiple requests (up to 15 if the request queue was empty) for a page to be promoted to the active LRU list to initiate the migration process. In the worst case, migrating one page may generate as many as 15 minor page faults.
@@ 3.2 Page Shadowing
To enable non-exclusive memory tiering, Matryoshka introduces a one-way page shadowing mechanism to enable a subset of pages resident in the performance tier to have a shadow copy in the capacity tier. Only pages promoted from the slow tier have shadow copies in the slow tier. Shadow copies are the original pages residing on the slow tier before they are unmapped in the page table and migrated to the fast tier. Shadow pages play a crucial role in minimizing the overhead of page migration during periods of memory pressure. Instead of swapping hot and cold pages between memory tiers, page shadowing enables efficient page demotion through page table remapping. This would eliminate half of the page migration overhead, i.e., page demotion, during memory thrashing.
@@
Indexing shadow pages. Inspired by the indexing of file-based data in the Linux page cache, Matryoshka builds an XArray for indexing shadow pages. An XArray is a radix-tree like, cache-efficient data structure that acts as a key-value store, mapping from the physical address of a fast tier page to the physical address of its shadow copy on the slow tier. Upon successfully completing a page migration, Matryoshka inserts the addresses of both the new and old pages into the XArray. Additionally, it adds a new shadow flag to the struct page of the new page, indicating that shadowing is on for this page.
@@
Shadow page management. The purpose of maintaining shadow pages is to assist page demotion. Fast or efficient page demotion is possible via page remapping if the master page, i.e., the one on the fast tier, is clean and consistent with the shadow copy. Otherwise, the shadow copy should be discarded. To track inconsistency between the master and shadow copies, Matryoshka sets the master page as read-only and a write to the page causes a page fault. To simplify system design and avoid additional cross-tier traffic, Matryoshka discards the shadow page if the master page is written.
@@
However, tracing updates to the master page poses a significant challenge. Page management in Linux relies heavily on the read-write permission of a page to perform various operations on the page, such as copy-on-write (CoW). While setting master pages as read-only effectively captures all writes, it may affect how these master pages are managed in the kernel. To address this issue, Matryoshka introduces a procedure called shadow page fault. It still designates all master pages as read-only but preserves the original read-write permission in an unused software bit on the page’s PTE (as shown in Figure 5). We refer to this software bit as shadow r/w. Upon a write to a master page, a page fault occurs. Unlike an ordinary page fault that handles write violation, the shadow page fault, which is invoked if the page’s shadow flag is set in its struct page, restores the read-write permission of the faulting page according to the shadow r/w bit and discards/frees the shadow page on the slow tier. The write may proceed once the shadow page fault returns and reinstates the page to be writable. For read-only pages, tracing shadow pages does not impose additional overhead; for writable pages, it requires one additional shadow page fault to restore their write permission.
@@
Reclaiming shadow pages. Non-exclusive memory tiering introduces space overhead due to the storage of shadow pages. If shadow pages are not timely reclaimed when the system is under memory pressure, applications may encounter out-of-memory (OOM) errors, which would not occur under exclusive memory tiering. There are two scenarios in which shadow pages should be reclaimed. First, the Linux kernel periodically checks the availability of free pages and if free memory falls below low_water_mark, kernel daemon kswapd is invoked to reclaim memory. Matryoshka instructs kswapd to prioritize the reclamation of shadow pages. Second, upon a memory allocation failure, Matryoshka also tries to free shadow pages. To avoid OOM errors, the number of freed shadow pages should exceed the number of requested pages. However, frequent memory allocation failures could negatively affect system performance. Matryoshka employs a straightforward heuristic to reclaim shadow pages, targeting 10 times the number of requested pages or until all shadow pages are freed. While excessive reclamation may have a negative impact on Matryoshka’s performance, it is crucial to prevent Out-of-Memory (OOM) errors. Experiments in Section 4 demonstrate the robustness of Matryoshka even under extreme circumstances.
3.3 Limitation
@@
Matryoshka relies on two rounds of TLB shootdown to effectively track updates to a migrating page during transactional page migration. When a page is used by multiple processes or mapped by multiple page tables, its migration involves multiple TLB shootdowns, per each mapping, that need to happen simultaneously. The overhead of handling multiple IPIs could outweigh the benefit of asynchronous page copy. Hence, Matryoshka deactivates transactional page migration for multi-mapped pages and resorts to the default synchronous page migration mechanism in Linux.
@@ 4 Evaluation
Figure 6: The three memory provisioning schemes used in evaluating tiered memory management. From bottom to top (with respect to fast memory): over-provisioning, approaching capacity, and under-provisioning.
This section presents a thorough evaluation of Matryoshka on its performance, overhead, and robustness. Most importantly, we aim at understanding tiered memory management through the comparison between Matryoshka and the existing representative approaches in three scenarios. We study two types of memory footprints: 1) resident set size (RSS) – the total size of memory occupied by a program; 2) working set size (WSS) – the amount of memory a program actively uses during execution. Since we focus on in-memory computing, WSS is typically smaller than RSS. Figure 6 shows the three scenarios we study to understand the tradeoff between migrating pages to the fast tier and accessing them directly from the slow tier. RSS determines the initial page placement while WSS decides the number of pages that should be migrated to the fast tier.
Platform A Platform B Platform C
@@
CPU 4th Gen Xeon Gold 4th Gen Xeon Platinum 2nd Gen Xeon Gold
DRAM 16 GB DDR5 16 GB DDR5 16 GB DDR4
CXL or PM Memory Agilex 7 16 GB DDR4 Agilex 7 16 GB DDR4 Optane 100 256 GB DDR-T
Table 1: The configurations of three testbeds.
Figure 7: Micro-benchmark: performance comparisons between TPP, Memtis, and Matryoshka with a WSS of 10 GB and a total of 20 GB of RSS data.
Figure 8: Micro-benchmark: performance comparisons between TPP, Memtis, and Matryoshka with a WSS of 13.5 GB and a total of 27 GB RSS.
@@
Small WSS (2.6M|1.5M)/((1.5M|1.48M) (3.3M|654K)/(1.63M|1.49M) (2|0)/(1|2) (0|0)/(0|0)
Medium WSS (6M|6M)(3.7M|3.9M) (6M|5.8M)/(3.7M|3.9M) (2.1M|2.3M)/(119K|3.35M) (2.1M|2M)/(116K|3.36M)
Large WSS (12.5M|12.3M)/(11.6M|11.5M) (13M|13M)/(12.3M|12M) (11.8M|11.6M)/(7.2M|11.3M) (11.9M|11.7M)/(7.2M|11.3M)
Table 2: The number of page promotions/demotions (read|write) during the initial in-process phase and the stable phase for TPP/Matryoshka. Note that, the number for Memtis is too small compared with TPP/Matryoshka, hence not listed.
Figure 9: Micro-benchmark: performance comparisons between TPP, Memtis, and Matryoshka with a WSS of 27 GB and the same 27 GB RSS.

### v2

@@ 2.3 Related Work
A long line of pioneering work has explored a wide range of tiered storage/memory systems, built upon SSDs and HDDs [266253, 8714805, chen2011hystor, 5749736, 10.5555/1855511.1855519, 6557143, 5581609, 6005391, 267006], DRAM and disks [270689, 270754, 269067, papagiannis2020optimizing], HBM and DRAMs [7056027, batman, 6493624], NUMA memory [autoNUMA, damon], PM and DRAM [288741, abulila2019flatflash, Adnan_22hpca, thermostat], local and far memory [Bergman_22ismm, 180194, 180727, 7266934, ruan2020aifm], DRAM and CXL memory [li2023pond, Maruf_23asplos_tpp, memtis], and multiple tiers [wu2021storage, Jonghyeon_21fast, Yan_19asplos_nimble, 288741, kwon2017strata]. We focus on tiered memory systems consisting of DRAM and the emerging byte-addressable memory devices, e.g., CXL memory and PM. Nomad also applies to other tiered memory systems such as HBM/DRAM and DRAM/PM.
@@
Lightweight memory access tracking. To mitigate software overhead associated with memory access tracking, Hotbox [Bergman_22ismm] employs two separate scanners for fast and slow tiers to scan the slow tier at a fixed rate while the fast tier at an adaptive rate, configurable based on the local memory pressure. Memtis [memtis] adjusts its PEBS-based sampling rate to ensure its overhead is under control (e.g., < 3%). TMTS [Padmapriya_23asplos] also adopts a periodic scanning mechanism to detect frequency along with hardware sampling to more timely detect newly hot pages. While these approaches balance scanning/sampling overhead and tracking accuracy, an “always-on” profiling component does not seem practical, especially for high-pressure workloads. Instead, thermostat [thermostat] samples a small fraction of pages, while DAMON [damon] monitors memory access at a coarser-grained granularity (i.e., region). Although both can effectively reduce the scanning overhead, coarse granularity leads to lower accuracy regarding page access patterns. On the other hand, to reduce the overhead associated with frequent hint page faults like AutoNUMA [autoNUMA], TPP [Maruf_23asplos_tpp] enables the page-fault based detection only for CXL memory (i.e., the slow tier) and tries to promote a page promptly via synchronous migration; prompt page promotion avoids subsequent page faults on the same page.
Inspired by existing lightweight tracking systems, such as Linux’s active and inactive lists and hint page faults, Nomad advances them by incorporating more recency information with no additional CPU overhead. Unlike hardware-assisted approaches [memtis, Padmapriya_23asplos, Loh2012ChallengesIH], Nomad does not require any additional hardware support.
@@
Page migration optimizations. To hide reclamation overhead from applications, TPP [Maruf_23asplos_tpp] decouples page allocation and reclamation; however, page migration remains in the critical path, incurring significant slowdowns. Nimble [Yan_19asplos_nimble] focuses on mitigating page migration overhead with new migration mechanisms, including transparent huge page migration and concurrent multi-page migration. Memtis [memtis] further moves page migration out of the critical path using a kernel thread to promote/demote pages in the background. TMTS [Padmapriya_23asplos] leverage a user/kernel collaborative approach to control page migration. In contrast, Nomad aims to achieve prompt, on-demand page migration while moving page migration off the critical path. It is orthogonal to and can benefit from existing page migration optimizations. The most related work is [concurrent], which leverages hardware support to pin data in caches, enabling access to pages during migration. Again, Nomad does not need additional hardware support.
3 Nomad Design and Implementation
@@
Nomad is a new page management mechanism for tiered memory that features non-exclusive memory tiering and transactional page migration. The goal of Nomad design is to enable the processor to freely and directly access pages from both fast and slow memory tiers and move the cost of page migration off the critical path of users’ data access. Note that Nomad does not make page migration decisions and relies on memory tracing the existing memory access tracking in the OS to determine hot and cold pagespage temperature. Furthermore, Nomad does not impact the initial memory allocation in the OS and assumes a standard page placement policy. Pages are allocated from the fast tier whenever possible and are placed in the slower tier only when there is an insufficient number of free pages in the fast tier, and or attempts to reclaim memory in the fast tier have failed. After the initial page placement, Nomad gradually migrates hot pages to the fast tier and cold pages to the slow tier. Nomad seeks to address two key issues: 1) how to minimize the cost of page migration? 2) how to minimize the number of migrations?
@@
Overview. Inspired by multi-level cache management in modern processors, which do not employ a purely inclusive or exclusive caching policy between tiers [Alian-micro21] to facilitate the sharing of or avoid the eviction of certain cache lines, Nomad embraces a non-exclusive memory tiering policy to prevent memory thrashing when under memory pressure. Unlike the existing page management schemes that move pages between tiers and require that a page is only present in one tier, Nomad instead copies pages from the slow tier to the fast tier and keeps a shadow copy of the migrated pages at the slow tier. The non-exclusive tiering policy maintains shadow copies only for pages that have been promoted to the fast tier, thereby not an inclusive policy. The advantage of the non-exclusive policy is that the demotion of clean, cold pages can be simplified to remapping the page table entry (PTE) without the need to actually copy the cold page to the slower tier.
@@
The building block of Nomad is a new transactional page migration (TPM) mechanism to reduce the cost of page migrations. Unlike the existing unmap-copy-remap 3-step page migration, TPM opportunistically copies a page without unmapping it from the page table. During the page copy, the page is not locked and can be accessed by a user program. After the copy is completed, TPM checks if the page has been dirtied during the copy. If not, TPM locks the page and remaps it in the PTE to the faster tier. Otherwise, the migration is aborted and will be tried at a later time. TPM not only minimizes the duration during which a page is inaccessible but also makes page migration asynchronous, thereby removing it from the critical path of users’ data access.
@@
Without loss of generality, we describe Nomad design in the context of Linux. We start with transactional page migration and then delve into page shadowing – an essential mechanism that enables non-exclusive memory tiering.
@@ 3.1 Transactional Page Migration
The motivation to develop TPM is to make page migration entirely asynchronous and decoupled from users’ access to the page. As discussed in Section 2.2, the current page migration in Linux is synchronous and on the critical path of users’ data access. For example, the default tiered memory management in Linux, TPP, attempts to migrate a page from the slow tier whenever a user program accesses the page. Since the page is in “protected” inaccessible mode, the access triggers a minor page fault, leading TPP to attempt the migration. The user program is blocked and makes no progress until the minor page fault is handled and the page is remapped to the fast tier, which can be a time-consuming process. Worse, if the migration fails, the OS remains in function migrate_pages and retries the aforementioned migration until it is successful or reaching a maximum of 10 attempts.
TPM decouples page migration from the critical path of user programs by making the migrating page accessible during the migrationprocessmigration. Therefore, users will access the migrating page from the slow tier before the migration is complete. While accessing a hot page from the slow tier may lead to sub-optimal memory performance, it avoids blocking user access due to the migration, thereby leading to superior user-perceived performance. Figure 3 shows the workflow of TPM. Before migration commences, TPM clears the protection bit of the page frame and adds the page to a migration pending queue. Since the page is no longer protected and it is not yet unmapped from the page table, following accesses to the page will not trigger additional page faults.
@@
Page migration is a complex procedure that involves memory tracing and updates to the page table for page remapping. The state-of-the-art page fault-based migration approaches, e.g., TPP in Linux [Maruf_23asplos_tpp], employs employ synchronous page migration, a mechanism in the Linux kernel for moving pages between NUMA nodes. In addition to the extended migration time affecting the critical path of user programs, this mechanism causes excessive page faults when integrated with the existing LRU-based memory tracing. TPP makes per-page migration decisions based on whether the page is on the active LRU list. Nevertheless, in Linux, memory tracing adds pages from the inactive to the active LRU list in batches of 15 requests  [^2] The 15 requests could be repeated requests for promoting the same page to the active LRU list, aiming to minimize the queue management overhead. Due to synchronous page migration, TPP may submit multiple requests (up to 15 if the request queue was is empty) for a page to be promoted to the active LRU list to initiate the migration process. In the worst case, migrating one page may generate as many as 15 minor page faults.
@@ 3.2 Page Shadowing
To enable non-exclusive memory tiering, Nomad introduces a one-way page shadowing mechanism to enable allow a subset of pages resident in the performance tier to have a shadow copy in the capacity tier. Only pages promoted from the slow tier have shadow copies in the slow tier. Shadow copies are the original pages residing on the slow tier before they are unmapped in the page table and migrated to the fast tier. Shadow pages play a crucial role in minimizing the overhead of page migration during periods of memory pressure. Instead of swapping hot and cold pages between memory tiers, page shadowing enables efficient page demotion through page table remapping. This would eliminate half of the page migration overhead, i.e., page demotion, during memory thrashing.
@@
Indexing shadow pages. Inspired by the indexing of file-based data in the Linux page cache, Nomad builds an XArray for indexing shadow pages. An XArray is a radix-tree like, cache-efficient data structure that acts as a key-value store, mapping from the physical address of a fast tier page to the physical address of its shadow copy on the slow tier. Upon successfully completing a page migration, Nomad inserts the addresses of both the new and old pages into the XArray. Additionally, it adds a new shadow flag to the struct page of the new page, indicating that shadowing is on for this page.
@@
Shadow page management. The purpose of maintaining shadow pages is to assist with page demotion. Fast or efficient page demotion is possible via page remapping if the master page, i.e., the one on the fast tier, is clean and consistent with the shadow copy. Otherwise, the shadow copy should be discarded. To track inconsistency between the master and shadow copies, Nomad sets the master page as read-only and a write to the page causes a page fault. To simplify system design and avoid additional cross-tier traffic, Nomad discards the shadow page if the master page is writtendirtied.
@@
However, tracing tracking updates to the master page poses a significant challenge. Page management in Linux relies heavily on the read-write permission of a page to perform various operations on the a page, such as copy-on-write (CoW). While setting master pages as read-only effectively captures all writes, it may affect how these master pages are managed in the kernel. To address this issue, Nomad introduces a procedure called shadow page fault. It still designates all master pages as read-only but preserves the original read-write permission in an unused software bit on the page’s PTE (as shown in Figure 5). We refer to this software bit as shadow r/w. Upon a write to a master page, a page fault occurs. Unlike an ordinary page fault that handles write violation, the shadow page fault, which is invoked if the page’s shadow flag is set in its struct page, restores the read-write permission of the faulting page according to the shadow r/w bit and discards/frees the shadow pageon the slow tier. The write may proceed once the shadow page fault returns and reinstates the page to be writable. For read-only pages, tracing tracking shadow pages does not impose additional overhead; for writable pages, it requires one additional shadow page fault to restore their write permission.
@@
Reclaiming shadow pages. Non-exclusive memory tiering introduces space overhead due to the storage of shadow pages. If shadow pages are not timely reclaimed when the system is under memory pressure, applications may encounter out-of-memory (OOM) errors, which would not occur under exclusive memory tiering. There are two scenarios in which shadow pages should be reclaimed. First, the Linux kernel periodically checks the availability of free pages and if free memory falls below low_water_mark, kernel daemon kswapd is invoked to reclaim memory. Nomad instructs kswapd to prioritize the reclamation of shadow pages. Second, upon a memory allocation failure, Nomad also tries to free shadow pages. To avoid OOM errors, the number of freed shadow pages should exceed the number of requested pages. However, frequent memory allocation failures could negatively affect system performance. Nomad employs a straightforward heuristic to reclaim shadow pages, targeting 10 times the number of requested pages or until all shadow pages are freed. While excessive reclamation may have a negative impact on Nomad’s performance, it is crucial to prevent Out-of-Memory (OOM) errors. Experiments in Section 4 demonstrate the robustness of Nomad even under extreme circumstances.
3.3 Limitations
@@
Nomad relies on two rounds of TLB shootdown to effectively track updates to a migrating page during transactional page migration. When a page is used by multiple processes or mapped by multiple page tables, its migration involves multiple TLB shootdowns, per each mapping, that need to happen simultaneously. The overhead of handling multiple IPIs could outweigh the benefit of asynchronous page copy. Hence, Nomad deactivates transactional page migration for multi-mapped pages and resorts to the default synchronous page migration mechanism in Linux. As high-latency TLB shootdowns based on IPIs continue to be a performance concern, modern processors, such as ARM, future AMD, and Intel x86 processors, are equipped with ISA extensions for faster broadcast-based [arm_tlb_broadcast, amd_tlb_broadcast] or micro-coded RPC-like [intelcpurpc] TLB shootdowns. These emerging lightweight TLB shootdown methods will greatly reduce the overhead of TLB coherence in tiered memory systems with expanded memory capacity. Nomad will also benefit from the emerging hardware and can be extended to scenarios where more intensive TLB shootdowns are necessary.
@@ 4 Evaluation
Figure 6: The three memory provisioning schemes used in the evaluation. From bottom to top concerning fast memory: over-provisioning, approaching capacity, and under-provisioning.
This section presents a thorough evaluation of Nomad, focusing on its performance, overhead, and robustness. Our primary goal is to understand tiered memory management by comparing Nomad with existing representative approaches to reveal the benefits and potential limitations of current page management appraches approaches for emerging tiered memory.
We analyze two types of memory footprints: 1) resident set size (RSS) – the total size of memory occupied by a program, and 2) working set size (WSS) – the amount of memory a program actively uses during execution. RSS determines the initial page placement, while WSS dictates the number of pages that should be migrated to the fast tier. Since we focus on in-memory computing, WSS is typically smaller than RSS. Figure LABEL:threescenarios 6 illustrates the three scenarios we study with the WSS size smaller than, close to, and larger than fast tier memory size.
Platform A Platform B Platform C Platform D
@@
CPU 4th Gen Xeon Gold 2.1GHz 4th Gen Xeon Platinum 3.5GHz 2nd Gen Xeon Gold 3.9GHz AMD Genoa 3.7GHz
Performance layer tier (DRAM) 16 GB DDR5 16 GB DDR5 16 GB DDR4 16GB DDR5
Capacity tier (CXL or PM Memory) Agilex 7 16 GB DDR4 Agilex 7 16 GB DDR4 Optane 100 256 GB DDR-T ×6 Micron CXL memory 256GB ×4
Performance layer tier read latency 316 cycles 226 cycles 249 cycles 391 cycles
Capacity layer tier read latency 854 cycles 737 cycles 1077 cycles 712 cycles
Performance tier bandwidth (GB/s) Single Thread / Peak performance Read: 12/31.45 Write: 20.8/28.5 Read: 12/31.2 Write: 22.3/23.67 Read: 12.57/116 Write: 8.67/85 Read: 37.8/270 Write: 89.8/272
Capacity tier bandwidth (GB/s) Single Thread/Peak performance Read: 4.5/21.7 Write: 20.7/21.3 Read: 4.45/22.3 Write: 22.3/22.4 Read: 4/40.1 Write: 8.1/13.6 Read: 20.25/83.2 Write: 57.7/84.3
Table 1: The configurations of four testbeds and performance characteristics of various memory devices.
@@
Small WSS (1.2M|1M)/(15.9K|134K)/ (1.16M|781K) (2.4M|2.2M)/(15.9K|140K)/ (2.7M|1.5M) (0|3.3K)/(7.7K|104K)/ (82|74) (424K|56K)/(0|104K)/ (48K|0)
Medium WSS (4M|6M)/(0|0)/ (1.6M|5M) (4.7M|6M)/(2|512)/ (2.5M|4.8M) (1.8M|3.2M)/(17.4K|0)/ (417K|1.6M) (1.9M|3.2M)/(16.9K|0)/ (293K|1.4M)
Large WSS (7M|5.9M)/(0|0)/ (4.5M|7M) (7.2M|6.5M)/(0|15)/ (4.1M|7.2M) (7.1M|5.2M)/(0|143K)/ (6.8M|8.8M) (7.1M|5.3M)/(0|143K)/ (6.8M|8.9M)
Table 2: The number of page promotions/demotions for read|write during the migration in progress and the stable phases for TPP/Memtis-Default/NOMAD. The data corresponds to Figure 7 for Platform platform A.